two blocks of masses m1=4.5 kg and m2=6.5 kg resting on a frictionless surface are connected by a light inextensible cord. A horizontal force F of 33N directed to the right is applied to the block with m1 and a horizontal force F of 11N directed to the left is applied to the block with m2. find the acceleration of the system and the tension in the rope.


1
Expert's answer
2021-11-15T13:04:40-0500

"33-T=4.5a"

"T-11=6.5a\to T=6.5a+11"


"33-6.5a-11=4.5a\to a=2\ (m/s^2)" . Answer


"T=6.5\cdot2+11=24\ (N)" . Answer